Use the ordinary annuity formula shown to the right to determine the accumulated amount in the annuity if ​$20 is invested semiannually for 10 years at 4.0​% compounded semiannually.
Answer #1

Answer:-

By using the Formula:
P=20 Semiannually
n=10*2=20
i=4.0%=0.04/2=0.02

Accumulated(Amount)=P[(1+i)^n-1]/i
=20[(1+0.02)^20-1]/0.02
=20[(1.02)^20-1]/0.02
=20(1.4859-1)/0.02
=20(.4859)/0.02
=9.718/0.02
   =$485.9
i.e. Accumulated amount=$485.9